OpenStudy (anonymous):

x^2-3x-10/2x^2-11x+5 / x^2-5x+6/2x^2-7x+3

OpenStudy (anonymous):

@austinL

OpenStudy (austinl):

\[\frac{\frac{(x^2-3x+10)}{(2x^2-11x+5)}}{\frac{(x^2-5x+6)}{(2x^2-7x+3)}}\]

OpenStudy (austinl):

Ok, you are dividing a fraction by a fraction, so what did I tell you that you should do in that case?

OpenStudy (anonymous):

I may have typed that wrong. it looks like this.

OpenStudy (anonymous):

but anyways, reciprocal

OpenStudy (austinl):

Exactly, you multiply the numerator by the reciprocal of the denominator. \[\frac{(x^2-3x-10)}{(2x^2-11x+5)} \times \frac{(2x^2-7x+3)}{(x^2-5x+6)}\]

OpenStudy (anonymous):

I don't know how to factor it

OpenStudy (austinl):

Ok, gimme a sec.

OpenStudy (austinl):

Lets start with the numerator of the first half. \[x^3-3x-10\]

OpenStudy (anonymous):

(x+5)(x-2)

OpenStudy (anonymous):

it's the 2x ones i don't get

OpenStudy (austinl):

FOIL that back out and see what you get? I don't believe that is correct.

OpenStudy (anonymous):

(x-5)(x+2)

OpenStudy (austinl):

Correct! Now, lets try one of those 2x ones.

OpenStudy (austinl):

What multiplied together is 2x^2?

OpenStudy (anonymous):

2x times x?

OpenStudy (austinl):

right so you would have this correct? \[(x-?)(2x-?)\]

OpenStudy (anonymous):

yeah

OpenStudy (austinl):

So, what times 2, plus what would give you 11?

OpenStudy (anonymous):

5 times 2 plus one?

OpenStudy (austinl):

Correct, so where would the 5 go, and where would the 1 go?

OpenStudy (anonymous):

so (x+5)(2x+1)

OpenStudy (austinl):

Correct, except they are negative because it is \[2x^2-11x+5\]

OpenStudy (austinl):

(x-5)(2x-1)

OpenStudy (anonymous):

so would the other one be (x-3)(2x-1) ?

OpenStudy (austinl):

Yes!

OpenStudy (austinl):

So what does that give us now? With it all factored out.

OpenStudy (anonymous):

so like this?

OpenStudy (austinl):

Correct!

OpenStudy (austinl):

\[\frac{(x+2)}{(x-2)}\]

OpenStudy (anonymous):

thanks so much! the first one was hard to read this is what it said

OpenStudy (austinl):

No problem. I was able to read it, it was just a little fuzzy and dark. You have another problem?

OpenStudy (anonymous):

I am having trouble with this one as well